OpenStudy (anonymous):

need formula for this question please! Hagan has a fixed rate mortgage of $534,000 at 5.4% for 20 years. The monthly payments are $3,643.22. How much will Hagan pay in interest after 20 years? $874,373 $562,836 $28,836 $340,373

OpenStudy (amistre64):

well, total cost of a loan amounts to what?

OpenStudy (amistre64):

its not about a formula, its about understanding whats actually working in the process. afterwards you can formulate to your hearts desire.

OpenStudy (anonymous):

Well then I don't understand, I keep getting the last one which is $340,373 I don't feel like its right

OpenStudy (amistre64):

then lets look at it this way do you agree that the total cost of a loan is what we pay for the loan amount and the interest? C = L + i

OpenStudy (amistre64):

this total cost is divided up into so many payments ... 12 times a year for 20 years C = 12(20)(P) since the cost doesnt change, we can equate the 2 setups and solve for interest 12(20)P = L + i

OpenStudy (anonymous):

but they put the interest rate in there? i'm lost now

OpenStudy (amistre64):

340 373 is fine

OpenStudy (amistre64):

the interest rate is just extra information .... you have to be able to parse thru the information to determine what is useless and what is important

OpenStudy (amistre64):

total payments is the total cost of the loan; which is the sum of its amount plus interest NP = L+i NP - L= i

OpenStudy (amistre64):

if we needed to determine what the monthly payments would be, then interest rate is valuable .... but they already give us this so its just redundant information

OpenStudy (anonymous):

okay soo my answer was correct right?

OpenStudy (amistre64):

12(20)(3643.22)-(534000) assuming my cranial math is good ... sure; lets dbl chk the wolf tho

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=12%2820%29%283643.22%29-%28534000%29 340 373 is fine

OpenStudy (anonymous):

okay thank you!

OpenStudy (amistre64):

yep